The demand function is Q = 200 - 5p, and the supply function is Q = -40 + 3p.determine the equilibrium price and quantity


1
Expert's answer
2021-11-18T10:21:17-0500

Solution:

At equilibrium: Qd = Qs

200 – 5p = -40 + 3p

200 + 40 = 3p + 5p

240 = 8p

P = 30

Equilibrium price = 30


Substitute in either the demand or supply function to derive equilibrium quantity:

Qd = 200 – 5p = 200 – 5(30) = 200 – 150 = 50

Qs = -40 + 3p = -40 + 3(30) = -40 + 90 = 50

Equilibrium quantity = 50 units